A business firms sells a good at the price of Rs 450.The firm has decided to reduce the price of good to Rs 350.Consequently, the quantity demanded for the good rose from 25,000 units to 35,000 units. Calculate the price elasticity of demand.
1
Expert's answer
2020-12-21T03:54:01-0500


Price elasticity of demand "= \frac{percentage change in demand}{percentage change in price}"


First of all, we have to find the midpoints to help us calculate percentage changes.


Demand midpoint "= \frac{25000+35000}{2} = 30,000"


Price midpoint "= \frac{350+450}{2} = 400"


So,

percentage change in demand "= \frac{35000-25000}{30000}\times100 = 33 \frac{1}{3}"


Percentage change in price "= \frac{350-450}{400}\times100 = -25"



Price elasticity of demand "= \frac{33.33}{-25} = -1\frac{1}{3}"